OpenStudy (anonymous):

Help me solve: sin(x + (π / 4)) – sin(x – (π / 4) = 1

rvc (rvc):

hey look at this diagram 

OpenStudy (anonymous):

Okay

rvc (rvc):

or else use the formula of \(\huge\color{red}{\tt sin(A-B)~and~sin(A+B)}\)

rvc (rvc):

do you know those formulas ?

OpenStudy (anonymous):

These are the addition and subtraction trig formulas for sine right?

rvc (rvc):

yes use them and expand the LHS

OpenStudy (anonymous):

LHS??? What's that

rvc (rvc):

Left Hand Side

OpenStudy (anonymous):

This is gonna take a while to type so hold on..

rvc (rvc):

Take your time :)

OpenStudy (anonymous):

sin ( a+b) = sin x cos (pi/4) + cos x sin (pi/4) sin (a - b) = -sin x cos (-pi/4) - cos x sin (pi/4)

OpenStudy (anonymous):

I already accounted for the "-sin" in the equation found in my question

rvc (rvc):

what sinx? ohh \(\Large(sin x cos (\pi/4) + cos x sin (\pi/4)\)\(\huge - \) \(\Large(sin x~ cos (-\pi/4) - cos x sin (\pi/4))\)

OpenStudy (anonymous):

So what's next?

rvc (rvc):

Saw something different i wrote?

rvc (rvc):

@Shorty1234 ?

OpenStudy (anonymous):

I'm confused...what are you talking about? @rvc

rvc (rvc):

Well the brackets are important when you open the brackets the sign changes you did not add the brackets

OpenStudy (anonymous):

Oh okay but we basically wrote the same thing except I didn't put the bracket in between "-" and "sin". So what's next?

rvc (rvc):

yeah so now open the brackets what do you get?

OpenStudy (anonymous):

What I wrote down the first time, right?

rvc (rvc):

be careful about the signs